Sales Mix and Break-Even Sales

Data related to the expected sales of laptops and tablets for Tech Products Inc. for the current year, which is typical of recent years, are as follows:

Products Unit Selling Price Unit Variable Cost Sales Mix
Laptops $1,400 $700 40%
Tablets 800 400 60%

The estimated fixed costs for the current year are $3,016,000.

Required:

1. Determine the estimated units of sales of the overall (total) product, E, necessary to reach the break-even point for the current year.
units

2. Based on the break-even sales (units) in part (1), determine the unit sales of both laptops and tablets for the current year.

Laptops units
Tablets units

3. Assume that the sales mix was 60% laptops and 40% tablets. Compare the breakeven point with that in part (1). Why is it so different?
units

The break-even point is in this scenario than in part (1) because the sales mix is toward the product with the higher of product.
Answer #2
Answer #1

1) Weighted average contribution margin per unit = (1400-700*40%)+(800-400*60%) = 520

Break even unit = 3016000/520 = 5800 Units

2. Based on the break-even sales (units) in part (1), determine the unit sales of both laptops and tablets for the current year.

Laptops (5800*40%) 2320 units
Tablets 3480 units

3) Weighted average contribution margin per unit = (1400-700*60%)+(800-400*40%) = 580

Break even unit = 3016000/580 = 5200 Units

Based on the break-even sales (units) in part (1), determine the unit sales of both laptops and tablets for the current year.

Laptops (5200*60%) 3120 units
Tablets 2080 units